OpenStudy (anonymous):

a zoo has 15 emperor penguins, which is 30% of the penguin population. how many penguins are there in all

OpenStudy (anonymous):

Transforming the words into an equation: \[pop \times 30 \%=pop \times 0.3 =15\]\[pop = \frac{15}{0.3}=\frac{3\times 5}{3 \times \frac{1}{10}}=50\]